### v3

@@ Abstract
Multi-agent trajectory data collected from domains such as team sports often suffer from missing values due to various factors. While many imputation methods have been proposed for spatiotemporal data, they are not well-suited for multi-agent sports scenarios where player movements are highly dynamic and inter-agent interactions continuously evolve. To address these challenges, we propose MIDAS (Multi-agent Imputer with Derivative-Accumulating Self-ensemble), a framework that imputes multi-agent trajectories with high accuracy and physical plausibility. It jointly predicts positions, velocities, and accelerations through a Set Transformer-based neural network and generates alternative estimates by recursively accumulating predicted velocity and acceleration values. These predictions are then combined using a learnable weighted ensemble to produce a final imputed trajectories. Experiments on three sports datasets demonstrate that MIDAS significantly outperforms existing baselines in both positional accuracy and physical plausibility. Lastly, we showcase use cases of MIDAS, such as approximating total distance and pass success probability, to highlight its applicability to practical downstream tasks that require complete tracking data.
@@ 1 Introduction
Many spatiotemporal domains such as transportation, robotics, surveillance, and sports handle multi-agent trajectory data. Although advances in computer vision and sensing technologies have facilitated large-scale trajectory data collection, acquiring complete trajectory data remains challenging due to various factors. Wearable devices such as GPS trackers and motion sensors may experience temporary signal loss or malfunction, and computer vision systems often miss agents due to occlusions or limitations of the camera’s field of view as instantiated in Fig. 1. This prevalence of missing values calls for the development of effective imputation techniques that can accurately reconstruct missing trajectories.
Though various imputation methods have been proposed for spatiotemporal data, applying them to multi-agent sports remains challenging due to their dynamic nature. In particular, while many of them [2, 5, 15, 16, 17, 23] have demonstrated their effectiveness in multi-sensor data such as including traffic flow or air quality datasets, they did not account for dynamic interactions between agents. In such fixed-sensor networks, spatial relationships are typically static and the underlying system structure remains constant over time. In contrast, multi-agent domains such as team sports involve players whose positions are continuously changing, requiring suitable architecture for explicitly modeling these dynamic inter-agent relationships while maintaining permutation-equivariance with respect to the agents. Furthermore, player motion is subject to biomechanical constraints that impose physical limits on their speed, acceleration, and directional changes. Therefore, imputed trajectories need to be not only accurate in terms of position but also physically plausible. However, due to the inherent uncertainty in player motion, existing imputation models [3, 18] even with dedicated architecture for multi-agent embedding tend to generate inaccurate or physically unrealistic trajectories, especially with limited training data or long missing intervals.
@@
Addressing these challenges, this paper proposes MIDAS (Multi-agent Imputer with Derivative-Accumulating Self-ensemble), a framework that imputes missing trajectories in multi-agent sports with high accuracy and physical plausibility by enforcing the physical constraints that real player trajectories should satisfy. First, a neural network equipped with Set Transformers [12] and player-wise bidirectional LSTMs [9] predicts the positions, velocities, and accelerations of missing players. In addition to this initial prediction (IP), it generates alternative estimates through derivative-accumulating prediction (DAP), which recursively accumulates the predicted velocity and acceleration values from the nearest observed positions for each missing segment in both directions. Finally, the weighted soft voting mechanism combines these three types of prediction, i.e. the IP and the forward/backward DAPs, by dynamically determining the weight of each prediction to produce the final imputed trajectories.
The proposed MIDAS overcomes the aforementioned challenges in three aspects. First, by employing the Set Transformer, it models dynamic inter-agent relationships while ensuring permutation-equivariance of agents. Second, by jointly predicting position, velocity, and acceleration and enforcing the physical consistency between them, it ensures that the imputed trajectories are not only accurate in terms of position, but also plausible and smooth in terms of velocity and acceleration. Third, through a learnable weighted ensemble, MIDAS improves the accuracy of the final imputed trajectories by addressing the physical inconsistency of IP and the error compounding problem of DAP. Experiments conducted on the three sports datasets demonstrate that MIDAS consistently outperforms existing baselines in terms of both positional accuracy and physical plausibility of imputed trajectories by a large margin.
In short, the main contributions of this paper include:
  1. 1.
    A Set Transformer-based imputation model specifically designed for multi-agent domains, which alone outperforms most existing baselines in sports trajectory imputation tasks.
  2. 2.
    The MIDAS mechanism, which simultaneously ensures accurate prediction of position, velocity, acceleration, and their physical consistency, resulting in producing both accurate and physically plausible trajectories.
  3. 3.
    Extensive experiments conducted on three benchmark datasets: soccer, basketball, and American football, demonstrating the generalizability and interoperability of the framework across different sports and data providers.
  4. 4.
    Real-world use cases of MIDAS, including the approximation of physical (e.g., total distance covered) and contextual (e.g., pass success probability) metrics based on imputed trajectories.
@@ 2 Related Work
Numerous methods have been proposed for time series and spatiotemporal data imputation. Early models such as BRITS [2], GRU-D [4], and MRNN [26] relied on sequential recurrence, but suffered from compounding errors due to their dependence on previous predictions. Non-autoregressive frameworks such as NAOMI [15], CSDI [23], and SAITS [7] mitigated this issue by enabling parallel imputation across time steps, improving robustness and efficiency. More recently, methods like TIDER [14] and TimesNet [24] were developed to capture the temporal patterns such as seasonality or local biases. Other recent approaches further leverage graph structures (GRIN [6], SPIN [16], and NRTSI [20]), information bottlenecks (TimeCIB [5]), or low-rank priors (ImputeFormer [17]) to enhance imputation performance. Although they perform well on static multi-sensor systems, they are not designed to handle dynamic interactions in multi-agent scenarios.
Meanwhile, several frameworks proposed dedicated architectures for multi-agent imputation to capture shifting spatial relationships in the domain. Notable examples include Graph Imputer [18] and TranSPORTmer [3], which employ graph networks [1] and set attention [12], respectively, to model dynamic player interactions while ensuring permutation equivariance. Interaction modeling is widely studied in future trajectory forecasting [8, 11, 13, 22, 25, 27], but existing methods are not directly applicable to missing trajectory imputation.
@@ 3.1 Neural Network-Based Initial Prediction
Since there is generally no inherent order among players in team sports, modeling their movements requires to ensure permutation-equivariance. That is, permuting the input order should not affect each player’s output, except for applying the same permutation in the output order. Following a recent study on the ball trajectory inference in team sports [10], we employ Set Transformer [12] to ensure the permutation-equivariance of outputs.
@@ 3.2 Derivative-Accumulating Prediction
In this section, we start from the fact that players’ acceleration values highly vary over time since it is directly related to their stochastic intents. In contrast, velocities are more correlated across neighboring time steps, and positions exhibit even stronger autocorrelation than their derivatives. This implies that accurately predicting acceleration values can lead to more stable and accurate position estimates, provided the physical relationships in Eq. (1) are maintained. However, since the model introduced in Section 3.1 does not enforce these relationships in its outputs $\{\hat{\mathbf{x}}^{k,i}_{1:T}\}_{k=1}^{K}$, accurate prediction of the derivatives does not necessarily lead to improved position accuracy.
@@ 4.1 Data Preparation
In the experiments, we independently train and evaluate models on three public datasets collected from popular team sports: soccer, basketball, and American football. The soccer dataset is provided by Metrica Sports [^1] https://github.com/metrica-sports/sample-data and contains tracking data for 22 players collected across three matches. For basketball, we use 100 matches among 631 matches of SportsVU NBA dataset [^2] https://github.com/linouk23/NBA-Player-Movements, which includes trajectories of 10 players per match. The American football dataset is from the Kaggle competition [^3] https://www.kaggle.com/competitions/nfl-big-data-bowl-2021 based on NFL’s Next Gen Stats. We adopt its preprocessed version [^4] https://github.com/lupalab/NRTSI/tree/main/codes_stochastic provided by NRTSI paper [20], which contains 9,543 five-second time series of six offensive players. The original sampling rates of the three datasets are 25Hz, 25Hz, and 10Hz, respectively, but we downsample all datasets to 10Hz for consistency. As model inputs, we use 200 frames (20 seconds) per sequence for soccer and basketball and 50 frames (5 seconds) for American football.
Table 1: Details on the Three Sports Datasets.
Soccer Basketball American Football
Split Matches Frames Matches Frames Matches Frames
Train. 2 70,119 70 1,843,329 - 425,000
Valid. 0.5 20,690 10 249,474 - 52,150
Test 0.5 21,424 20 529,764 - -
@@ 4.2 Missing Scenarios
Figure 3: Visualization of example masking matrices for three different missing scenarios.
Fig. 3 illustrates examples of masking matrices from the soccer dataset used in each missing scenario. Since our task is to impute trajectories given observed data before and after the missing intervals, we ensure that the first and last five frames of each sequence are always observable, following previous work [15, 18, 20]. During training, we apply a dynamic missing rate ranging from 0.1 to 0.9, while testing is conducted with a fixed missing rate of 0.5 to encourage the model to generalize across various missing rate scenarios. Additional evaluations with different missing rates are presented as an the ablation study in Section 4.5.
@@ 4.3 Baseline Models and Evaluation Metrics
Table 2: Performance of imputation models on different datasets and scenarios.
@@
4.4 Main Experimental Results
The resulting Table 2 shows that the proposed MIDAS consistently outperforms other baselines in both positional accuracy (PE) and physical plausibility (SCE). Linear interpolation (LI) and cubic spline (CS) offer simple yet competitive results in some scenarios. Notably, deep learning baselines such as BRITS, NAOMI, and CSDI, which are not designed for multi-agent domains, often fail to exceed these naive baselines. Similarly, Graph Imputer (GI) performs worse than LI and CS in many cases despite its dedicated architecture for sports contexts, as previously reported in its original article [18].
By contrast, our model for initial prediction (IP) already surpasses most baselines, demonstrating its strength in modeling dynamic multi-agent interactions. Furthermore, comparing IP and MIDAS highlights the effectiveness of our self-ensemble mechanism. By effectively leveraging the complementary strengths of intial predictions and alternative derivative-accumulating predictions, it leads to superior performance even compared to state-of-the-art methods such as ImputeFormer (IF) in most scenarios.
@@ 4.4 Main Experimental Results
Figure 4: Comparison of ground truth and imputed trajectories for the uniform missing scenario in basketball. The solid lines indicate the ground truth trajectories, while the dashed ones show imputed trajectories resulting from different methods. Circles indicate the players’ locations at the last frame, and the annotated values represent average position errors over the missing interval.
Meanwhile, an interesting observation emerges in the basketball dataset, where IP and IF show performance comparable to or slightly better than MIDAS in some cases. We attribute this to the larger amount of training data available (70 games), which may reduce the relative advantage of MIDAS. In contrast, for soccer and American football, where training data are more limited, MIDAS’s derivative-based self-ensemble mechanism demonstrates its data efficiency by significantly outperforming other methods. This data efficiency is particularly important in real-world sports applications, where acquiring large amounts of complete tracking data is often difficult due to the competitive and commercial nature of professional sports.
@@ 4.5 Ablation Studies
In addition to the main experiments, we carried out an ablation study to examine the impact of each component of MIDAS. We also conducted a comparative analysis on the performance of the framework under different sequence lengths and missing rates. For simplicity, both experiments were conducted only on the broadcasting camera scenario in the soccer dataset.
Impact of the Derivatives as Features
To empirically justify the use of derivative features in IP and DAP, we compared the performance of MIDAS on the broadcasting camera scenario with those trained without accelerations (i.e., using positions and velocities) and even without velocities (i.e., only using positions) of observed trajectories. Furthermore, we configured IP predict only positions and velocities, and DAP estimate positions only based on predicted velocities. We then compared these “velocity-only” DAP models with their counterparts that also incorporate predicted accelerations as in Eq. (7).
Table 3: Position errors of the components of MIDAS in Eq. (11) trained using different subsets of features.
@@
As shown Table 3, providing velocities and accelerations as input features for IP clearly improves the imputation performance. We attribute this to the neural network architecture for IP, where Set Transformers independently encode multi-agent contexts at each time step via Set Transformers, and player-wise Bi-LSTMs link temporal information. Providing derivatives as input allows the Set Transformers to utilize information of adjoining time steps, resulting in more comprehensive context embeddings compared to using positions alone.
@@
Results on Different Sequence Lengths and Missing Rates
In addition, we compare the model’s performance under different sequence lengths and missing rates. Agreeing with our intuition, Table 4 shows that the model becomes less accurate as either the sequence length or the missing rate increases. In particular, the position error of the agent-wise missing scenario becomes much larger for a longer sequence, even with a small missing rate. This is because some players may miss most frames in the process of randomly distributing the total number of missing frames to individuals. Accordingly, in a longer sequence, the model has to predict the players’ stochastic movements for tens of seconds, resulting in a large error due to its inherent difficulty. Meanwhile, in the broadcasting camera scenario, the maximum length of a missing segment does not significantly change according to the sequence length. This results in robust performance against varying sequence lengths, implying that our framework can be reliably applied to real-world broadcasting data.
Table 4: Position errors of the MIDAS trained for and applied to scenarios with different sequence lengths and missing rates. We report the missing rate of the camera scenario alongside since it varies according to the sequence length.
SL \,MR 0.3 0.5 0.7 Camera (MR)
50 0.0691 0.1158 0.1552 0.0841 (0.4569)
100 0.3561 0.5948 0.8047 0.4622 (0.5138)
200 1.2805 2.0491 2.5556 1.2296 (0.5375)
300 2.3129 3.2746 4.0821 1.5993 (0.5454)
600 4.2597 5.9329 7.2485 2.1410 (0.5540)
4.6 Computational Complexity Analysis
In this section, we analyze the computational complexity of MIDAS. The analysis considers three key parameters: the number of time steps $T$, the hidden dimension $m$, and the number of players $K$. It takes $O(TK^{2}m)$ for Set Transformer used in the initial prediction, $O(TKm^{2})$ for Bi-LSTMs employed in the initial prediction and weighted ensemble. In total, $O(TK^{2}m+TKm^{2})\approx O(TKm^{2})$ as $m\gg K$. Since Bi-LSTM adopts a recurrent structure, the entire computation process proceeds over $O(T)$ steps, proportional to the length of the sequence.

### v4

@@ Abstract
Multi-agent trajectory data collected from domains such as team sports often suffer from missing values due to various factors. While many imputation methods have been proposed for spatiotemporal data, they are ill-suited for multi-agent sports, where player movements are highly dynamic and interactions evolve over time. To address these challenges, we propose MIDAS (Multi-agent Imputer with Derivative-Accumulating Self-ensemble), a data-efficient framework that imputes multi-agent trajectories with high accuracy and physical plausibility. It jointly predicts positions, velocities, and accelerations via a Set Transformer-based neural network and refines them by recursively accumulating predicted velocity and acceleration values. These predictions are then combined using a learnable weighted ensemble to produce final imputed trajectories. Experiments on three sports datasets show that MIDAS significantly outperforms existing baselines, with particularly large margins in limited-data settings. We also demonstrate its utility in downstream tasks such as estimating total distance and pass success probability. The source code is available at https://github.com/gkswns95/midas.git.
@@ 1 Introduction
Many spatiotemporal domains, such as transportation, robotics, surveillance, and sports, handle multi-agent trajectory data. While advances in computer vision and sensing technologies have facilitated large-scale trajectory data collection, acquiring complete data remains challenging due to various factors such as signal loss in wearable devices and limitations of the camera’s field of view (Fig. 1). This prevalence of missing values calls for the development of effective imputation techniques that can accurately reconstruct missing trajectories.
@@
Though various imputation methods have been proposed for spatiotemporal data, applying them to multi-agent sports remains challenging due to their dynamic nature. In particular, while many of them [1, 5, 16, 17, 18, 24] have demonstrated their effectiveness in multi-sensor data, including traffic flow or air quality datasets, they do not account for dynamic interactions between agents. In such fixed-sensor networks, spatial relationships typically remain constant over time. In contrast, multi-agent domains such as team sports involve players with continuously changing positions, requiring a suitable architecture for explicitly modeling these dynamic inter-agent relationships while maintaining permutation-equivariance with respect to the agents.
Furthermore, player motion is governed by biomechanical constraints that impose physical limits on speed, acceleration, and directional change. To accurately model such constraints and generate physically plausible trajectories, imputation models must learn nuanced patterns in movement dynamics. This typically requires a large volume of high-quality training data. However, this poses a significant limitation in the sports domain, where player tracking data is often treated as confidential due to the competitive nature of professional leagues.
@@
Addressing these challenges, this paper proposes MIDAS (Multi-agent Imputer with Derivative-Accumulating Self-ensemble), a framework for multi-agent sports that imputes missing trajectories with high accuracy and data efficiency by explicitly enforcing the physical constraints that real player trajectories should satisfy. First, a neural network equipped with Set Transformers [13] and player-wise bidirectional LSTMs [9] predicts the positions, velocities, and accelerations of missing players. In addition to this initial prediction (IP), it generates alternative estimates through derivative-accumulating prediction (DAP), which recursively accumulates the predicted velocity and acceleration values from the nearest observed positions in both forward and backward directions. Finally, the learnable soft voting mechanism combines these three predictions, namely IP and forward/backward DAPs, to produce the final imputed trajectories.
The proposed MIDAS overcomes the aforementioned challenges in two aspects. First, by employing the Set Transformer, it models dynamic inter-agent relationships while ensuring permutation-equivariance with respect to agents. Second, MIDAS jointly predicts positions, velocities, and accelerations, and integrates them through a derivative-accumulating self-ensemble mechanism that explicitly enforces their physical consistency. This design enables the model to capture the underlying patterns of player motion, leading to improved data efficiency.
Experiments across three sports datasets demonstrate that MIDAS consistently outperforms existing baselines in terms of both positional accuracy and physical plausibility of imputed trajectories, with particularly large margins in low-data settings. In addition, we showcase real-world applications of MIDAS, including the approximation of physical (e.g., total distance covered) and contextual (e.g., pass success probability) metrics based on imputed trajectories.
@@ 2 Related Work
Numerous methods have been proposed for time series and spatiotemporal data imputation. Early models such as BRITS [1], GRU-D [4], and MRNN [28] relied on recurrent models, but suffered from compounding errors due to dependence on their previous predictions. Non-autoregressive frameworks such as NAOMI [16], CSDI [24], and SAITS [7] mitigated this issue by enabling parallel imputation across time steps, improving robustness and efficiency. More recently, methods like TIDER [15] and TimesNet [25] were developed to capture temporal patterns such as seasonality or local biases. Other recent approaches further leverage graph structures (GRIN [6], SPIN [17], and NRTSI [21]), information bottlenecks (TimeCIB [5]), or low-rank priors (ImputeFormer [18]) to enhance imputation performance. Although they perform well on static multi-sensor systems, they are not designed to handle dynamic interactions in multi-agent scenarios.
Meanwhile, several frameworks have proposed dedicated architectures for multi-agent imputation to capture shifting spatial relationships of agents. Notable examples include Graph Imputer [19], GC-VRNN [26], and TranSPORTmer [2], which employ dynamic graph or set attention [13] architectures to model time-varying player interactions. More recently, methods tailored for sports data with advanced architectures have been explored, such as U2Diff [3] and Event2Tracking [10]. However, the former suffers from high computational costs and low data efficiency inherent to diffusion models, while the latter requires manually annotated event data, which are not always available. Interaction modeling has been widely studied in future trajectory forecasting [8, 12, 14, 23, 27, 29], but these methods are not directly applicable to missing trajectory imputation.
@@ 3.1 Neural Network-Based Initial Prediction
Since there is generally no inherent order among players in team sports, modeling their movements requires ensuring permutation-equivariance. That is, permuting the input order should not affect each player’s output, except for applying the same permutation in the output order. Following a recent study on the ball trajectory inference in team sports [11], we employ Set Transformer [13] to ensure the permutation-equivariance of outputs.
@@ 3.2 Derivative-Accumulating Prediction
In this section, we start from the fact that players’ acceleration values highly vary over time since they are directly related to their stochastic intents. In contrast, velocities are more correlated across neighboring time steps, and positions exhibit even stronger autocorrelation than their derivatives. This implies that accurately predicting acceleration values can lead to more stable and accurate position estimates, provided the physical relationships in Eq. (1) are maintained. However, since the model introduced in Section 3.1 does not enforce these relationships in its outputs $\{\hat{\mathbf{x}}^{k,i}_{1:T}\}_{k=1}^{K}$, accurate prediction of the derivatives does not necessarily lead to improved position accuracy.
@@ 3.3 Weighted Ensemble of Multiple Predictions
Fig. 3 illustrates how MIDAS dynamically adjusts ensemble weights depending on the characteristics of the missing segment. For Player 2, who has a short and stable missing trajectory, the model assigns negligible weight to IP, relying more on DAPs for imputation. In contrast, Player 3’s longer and more variable missing trajectory leads to higher IP weights, especially in the middle where DAP errors may accumulate. This highlights MIDAS’s ability to improve the final prediction by adaptively combining its components.
Figure 3: Ensemble weight examples for individual imputed trajectories, where circles indicate each player’s initial position.
@@ 4.1 Data Preparation
In the experiments, we independently trained and evaluated models on three public datasets collected from popular team sports: soccer, basketball, and American football. The soccer dataset is provided by Metrica Sports [^1] https://github.com/metrica-sports/sample-data and contains tracking data for 22 players collected across three matches. For basketball, we use the first 100 matches of SportsVU NBA dataset [^2] https://github.com/linouk23/NBA-Player-Movements, containing trajectories of 10 players per match. The American football dataset is from the Kaggle competition [^3] https://www.kaggle.com/competitions/nfl-big-data-bowl-2021 and is based on NFL’s Next Gen Stats. We adopt its preprocessed version [^4] https://github.com/lupalab/NRTSI/tree/main/codes_stochastic provided by the NRTSI paper [21], which contains 9,543 five-second time series of six offensive players. The original sampling rates of the three datasets are 25Hz, 25Hz, and 10Hz, respectively, but we downsample all datasets to 10Hz for consistency. As model inputs, we use 200 frames (20 seconds) per sequence for soccer and basketball and 50 frames (5 seconds) for American football.
@@ 4.2 Missing Scenarios
Fig. 4 illustrates examples of masking matrices from the soccer dataset used in each missing scenario. Since our task is to impute trajectories given observed data before and after the missing intervals, we ensure that the first and last five frames of each sequence are always observable, following previous work [16, 19, 21]. During training, we apply a dynamic missing rate ranging from 0.1 to 0.9, while testing is conducted with a fixed missing rate of 0.5 to encourage the model to generalize across various missing rate scenarios. Additional evaluations with different missing rates are presented as the ablation study in Section 4.5.
Table 1: Details on the three sports datasets.
Soccer Basketball A. Football
Split Matches Frames Matches Frames Matches Frames
Training 2 65,014 70 1,621,835 425,000
Validation 0.5 20,104 10 216,118 52,150
Test 0.5 21,242 20 468,885
Figure 4: Masking matrix examples for three different missing scenarios.
@@ 4.3 Baseline Models and Evaluation Metrics
4.4 Main Experimental Results
The resulting Table 2 shows that the proposed MIDAS consistently outperforms other baselines in both positional accuracy (PE) and physical plausibility (SCE). Linear interpolation (LI) and cubic spline (CS) offer simple yet competitive results in some scenarios. Notably, deep learning baselines such as BRITS, NAOMI, and CSDI, which are not designed for multi-agent domains, often fail to exceed these naive baselines. Similarly, Graph Imputer (GI) performs worse than LI and CS in many cases, as previously reported in its original article [19].
By contrast, our model for initial prediction (IP) already surpasses most baselines, demonstrating its strength in modeling dynamic multi-agent interactions. Furthermore, comparing IP and MIDAS highlights the effectiveness of our self-ensemble mechanism. By effectively leveraging the complementary strengths of initial predictions and alternative derivative-accumulating predictions, it leads to superior performance even compared to state-of-the-art methods such as ImputeFormer (IF) in most scenarios.
Table 2: Performance of imputation methods on different datasets and scenarios.
@@ 4.4 Main Experimental Results
Table 3: Performance of the top-3 methods on the basketball dataset when trained with full data (70 games) versus limited data (3 games, 71,428 frames).
Limited Data Full Data
Scenario Metric IF IP MIDAS IF IP MIDAS
Uniform PE 1.6741 1.1868 1.1438 1.3622 0.9801 0.9727
SCE 0.0628 0.1483 0.0493 0.0531 0.0432 0.0438
Agent-wise PE 1.8876 1.6414 1.5994 1.3345 1.3832 1.3862
SCE 0.0645 0.0452 0.0439 0.0485 0.0373 0.0381
@@
Meanwhile, an interesting observation emerges in the basketball dataset, where IP and IF show performance comparable to or slightly better than MIDAS in some cases. We attribute this to the larger amount of training data available (70 games), which may reduce the relative advantage of MIDAS. In contrast, for soccer and American football, where training data are more limited, MIDAS’s derivative-based self-ensemble mechanism demonstrates its data efficiency by significantly outperforming other methods.
To investigate whether this data efficiency holds for the basketball dataset, we measure model performance on an additional setting where models are trained using only 3 games (comparable in scale to the training data used for the soccer dataset). As shown in Table 3, all models experience a performance drop when trained on limited data. However, the degradation is much less severe for MIDAS, which eventually outperforms both IF and IP under this setting. This observation underscores the practical utility of MIDAS in real-world sports analytics, where obtaining large amounts of complete tracking data is often difficult due to the competitive and commercial nature of professional sports.
@@
Figure 5: Ground truth (solid) and imputed trajectories (dashed) for a uniform missing scenario in basketball. Circles indicate the players’ initial positions, and the annotated values represent average position errors over the missing interval.
@@ 4.5 Ablation Studies
In addition to the main experiments, we conducted two ablation studies to further investigate the model’s behavior across varying difficulties and the impact of its features and components on the performance.
Model Behavior Across Varying Difficulties.
To better understand how MIDAS responds to varying levels of difficulty, we conduct an ablation study analyzing the model’s behavior across different missing lengths. For the agent-wise missing scenario in the soccer and basketball datasets, we divide trajectory segments in the test set into three groups based on missing length.
[…]
Sports Category Missing Frames $\hat{\mathbf{p}}^{k,i}_{t}$ ($\hat{\lambda}^{k,i}_{t}$) $\hat{\mathbf{p}}^{k,f}_{t}$ ($\hat{\lambda}^{k,f}_{t}$) $\hat{\mathbf{p}}^{k,b}_{t}$ ($\hat{\lambda}^{k,b}_{t}$) $\hat{\mathbf{p}}^{k}_{t}$
Soccer Short 033.30±15.72plus-or-minus33.3015.7233.30$\pm$ 15.7233.30 ± 15.72 0.1379 (0.0001) 0.0742 (0.6501) 0.0783 (0.3498) 0.0504
Medium 090.42±19.17plus-or-minus90.4219.1790.42$\pm$ 19.1790.42 ± 19.17 0.7939 (0.0003) 0.7492 (0.6110) 0.7624 (0.3887) 0.7004
Long $173.41\pm 21.70$ 2.7359 (0.0002) 2.7017 (0.5034) 2.7186 (0.4964) 2.6082
Basketball Short 033.86±16.11plus-or-minus33.8616.1133.86$\pm$ 16.1133.86 ± 16.11 0.0444 (0.1596) 0.0424 (0.5216) 0.0426 (0.3187) 0.0376
Medium 090.65±19.02plus-or-minus90.6519.0290.65$\pm$ 19.0290.65 ± 19.02 0.5395 (0.1783) 0.5483 (0.4809) 0.5482 (0.3406) 0.5312
Long $172.60\pm 21.20$ 1.8104 (0.1801) 1.8288 (0.4313) 1.8258 (0.3985) 1.8005
Table 5: Position errors of each MIDAS component in Eq. (11), when trained using different subsets of features.
@@
As shown in Table 4, the position error increases with longer missing segments, and the advantage of DAP over IP diminishes. When the missing length is short, the DAP components outperform the initial prediction (IP) by a wide margin, resulting in substantial performance gains through the ensemble. However, as the missing length increases, the uncertainty of player motion makes DAP less reliable, reducing its advantage. This trend is reflected in the ensemble weights as well: $\hat{\lambda}^{k,i}_{t}$ increases and $\hat{\lambda}^{k,f}_{t}$ decreases as the missing length grows.
In addition, the advantage of DAP over IP varies by sport. In soccer, DAP maintains its superiority over IP even for long segments, and the IP component receives almost zero weight in the ensemble, indicating that MIDAS relies entirely on DAP. In contrast, in basketball, IP surpasses DAP for medium and long segments. This is reflected in noticeably higher $\hat{\lambda}^{k,i}_{t}$, suggesting its greater influence in ensemble prediction. This difference likely stems from the nature of basketball, where the variability of player motion is higher relative to the small court size, making long-range accumulation of derivatives more error-prone.
Impact of Derivative Features.
To empirically justify the use of derivative features in IP and DAP, we compared the performance of MIDAS on the broadcasting camera scenario in soccer with those trained without accelerations (i.e., using positions and velocities) and even without velocities (i.e., only using positions) of observed trajectories. Furthermore, we configured IP to predict only positions and velocities, and DAP to estimate positions only based on predicted velocities. We then compared these “velocity-only” DAP models with their counterparts that also incorporate predicted accelerations as in Eq. (7).
Table 5 shows that providing velocities and accelerations as input features for IP clearly improves the imputation performance. We attribute this to the neural network architecture for IP, where Set Transformers independently encode multi-agent contexts at each time step via Set Transformers, and player-wise Bi-LSTMs link temporal information. Providing derivatives as input allows the Set Transformers to utilize information of adjoining time steps, resulting in more comprehensive context embeddings compared to using positions alone.
@@
4.6 Time Complexity Analysis
While achieving high imputation accuracy is important, real-world deployment requires efficient processing without excessive computational overhead. To assess the feasibility of real-time use, we analyze both the theoretical and empirical time complexity of the proposed MIDAS framework.
Table 6: Empirical analysis of time complexity on soccer and basketball data. Each row corresponds to an imputation method, with the number of parameters, average inference time per 20-second window, and total processing time for half a soccer match and a full basketball match in the test set described in Table 1.
Soccer Basketball
Method Params Inference Total Params Inference Total
[…]
From a theoretical standpoint, we consider three key parameters: the number of time steps $T$, the hidden dimension $m$, and the number of players $K$. It takes $O(TK^{2}m)$ for Set Transformer used in the initial prediction, $O(TKm^{2})$ for Bi-LSTMs employed in the initial prediction and weighted ensemble. In total, $O(TK^{2}m+TKm^{2})\approx O(TKm^{2})$ as $m\gg K$. Since Bi-LSTM adopts a recurrent structure, the entire computation process proceeds over $O(T)$ steps, proportional to the length of the sequence.
In addition to this analysis, we empirically evaluated the inference time of MIDAS and several baseline models using an NVIDIA TITAN RTX GPU with 24GB memory. As summarized in Table 6, although MIDAS is slightly slower than ImputeFormer [18], it can process an entire 40-minute game (half-match for soccer and full game for basketball) within 10 seconds. This demonstrates that MIDAS operates well within real-time constraints, enabling its potential use in practical scenarios such as live match analysis or broadcasting augmentation.